OpenStudy (marmar10):

Help! how fast does a .4kg ball hit the ground when it is dropped from a height of 50m?

OpenStudy (perl):

you can use the equation y = 1/2 * g * t^2 + y0 First find the time it takes to hit the ground, then plug this time into the velocity equation dy/dt = g*t

OpenStudy (anonymous):

Follow @perl One of the most fundamental ideas in physics is that under the influence of gravity the mass of an object does not affect the acceleration of the object. All objects on Earth accelerate at 9.8 m/s^2 The mass here is irrelevant in this question.

OpenStudy (anonymous):

I would just use KE=1/2mv^2 and Ug=mgh ... Since all the potential gravitational (Ug) energy will be transfered to Kinetic energy (KE) The equation will come out as... \[v=\sqrt{2gh}\]

OpenStudy (marmar10):

thanks :)